OpenStudy (anonymous):

George is six inches taller that his son, Reese. Their combined height is 134 inches. George is _____ " inches tall Reeses is _____" inches tall Please help and explain :) Medals given!

OpenStudy (nikato):

U should set up two equations. Let G=George and R=Reese

OpenStudy (nikato):

So how would u write George is 6 inches taller than Reese as an equation?

OpenStudy (anonymous):

George is 70 inches and Reese is 67 inches. Amen may lord help us all. #Mindblown #Swag #TheAnswer #SoManyhashtags #Thisain'ttwitter

OpenStudy (anonymous):

134 = G + R G = R+ 6 plug this in for G above 134 = (R+6) + R 134 = 2R + 6 134-6 = 2R R = (134-6)/2=64 Now solve for G G = 64 + 6 = 70 Now check 134 = 70 + 64

OpenStudy (anonymous):

WAit actually rees is 64 mb.

OpenStudy (anonymous):

inches*

OpenStudy (anonymous):

Thank you!!!!!!!!!!!!!!!!:)

OpenStudy (anonymous):

np here is a medal.

OpenStudy (anonymous):

Do you understand or did you just want the answer?

OpenStudy (nikato):

No, George is 64

OpenStudy (anonymous):

I understood it..

OpenStudy (anonymous):

Okay good!

OpenStudy (anonymous):

so reese is 70?

OpenStudy (anonymous):

wait.. I confused

OpenStudy (anonymous):

noooo lets, say Reese is x x + (x + 6) = 134 2x + 6 = 134 2x = 128 x = 64 inches (height of Reese) 64 + 6 = 70 inches (height of George)

OpenStudy (anonymous):

@nikato right?

OpenStudy (nikato):

Nvm, George is 70

OpenStudy (nikato):

My G looked like a 6

OpenStudy (anonymous):

Yea I said 70.